# Scheduling problem of aircrafts on a same runway and dual runways

### v3

@@ VI Scheduling problem of landing and takeoff aircraft on dual runways
If one aircraft $Tcf_{i}$ lands and one aircraft $Tcf_{j}$ takes off at the same time, it is assumed by default that aircraft $Tcf_{i}$ is ahead of aircraft $Tcf_{j}$, i.e., $S_{i}\leq S_{j}$ the aircraft $Tcf_{i}$ is the leading aircraft and $Tcf_{j}$ is the trailing aircraft, and it is said that the aircraft $Tcf_{j}$ is relevant to the aircraft $Tcf_{i}$.
@@
(2) Suppose that aircraft $Tcf_{i}$ and $Tcf_{i+1}$ are both landing (takeoff) aircraft. If aircraft $Tcf_{i+1}$ is relevant to some aircraft $Tcf_{k}$, $k=i$.
@@
(2) If the optimization problem (6) can be solved, the aircraft $Tcf_{2}$ is relevant to the aircraft $Tcf_{1}$ and for each $i\in\{3,4,\cdots,n\}$, the aircraft $Tcf_{i}$ is relevant to an aircraft $Tcf_{k}$ for $k\in\{Tcf_{i-5},\cdots,Tcf_{i-2},Tcf_{i-1}\}$.
@@
In Theorem 13, the aircraft $Tcf_{k+1}$ is actually a semi-resident-point aircraft. From Theorem 13, the relative time increment is $T_{0}$ when the exchange of aircraft $Tcf_{k}$ and $Tcf_{k+1}$ are exchanged, which is a significant property for the algorithm proposed later.
@@
From Theorem 16, it can be observed that the unequal separation times caused possibly by the existence of breakpoint aircraft and resident-point aircraft might result in mismatches between the landing sequence $\Phi_{0}$ and the takeoff sequence $\Phi_{1}$, but can be restored after $1$ to $4$ aircraft in the absence of the influence of other breakpoint aircraft and resident-point aircraft. Moreover, it should be noted here that the forms of the minimum separation times additionally with a sequence such as $Y_{j(j+1)}(\Phi_{1})$ and $Y_{(j+1)(j+2)}(\Phi_{1})$ are used to distinguish different separation times between different sequences.
@@
(1) Suppose that $Tcf_{i_{1}-1}\in\phi_{0}$, $Tcf_{i_{1}}\in\phi_{1}$, for the sequence $\phi_{0}$, there is a path from its last aircraft $Tcf_{i_{1}-1}$ to its first aircraft, and for the sequence $\phi_{1}$, there is a path from its second last aircraft to its first aircraft. If $Tcf_{i_{1}}$ is only relevant to $Tcf_{i_{1}-1}$, it is said that the sequence $\phi$ is a T-block subsequence of $\Phi$, and $(|\phi_{1}|-1)/(|\phi_{0}|-1)$ and $S_{i_{1}}(\phi)-S_{i_{2}}(\phi)-Y_{i_{2}i_{1}}$ are the length and the takeoff time increment of T-block subsequence, where $Tcf_{i_{2}}$ is the second last takeoff aircraft in $\phi$. Let $\langle Tcf_{k_{0}},Tcf_{k_{1}},\cdots,Tcf_{k_{s}}\rangle$ denote the takeoff sequence in the T-block subsequence $\phi$. Consider a subsequence $\bar{\phi}$ such that the landing times of the landing aircraft in $\bar{\phi}$ are consistent with those in $\phi$ and the takeoff times of the takeoff aircraft satisfy that $S_{k_{0}}(\bar{\phi})-S_{k_{0}}(\phi)=S_{k_{1}}(\bar{\phi})-S_{k_{1}}(\phi)=\cdots=S_{k_{s-1}}(\bar{\phi})-S_{k_{s-1}}(\phi)=\mu$ and $S_{k_{s}}(\bar{\phi})=S_{k_{s}}(\phi)$ for some constant $\mu\geq 0$. Suppose that for a given constant $\mu_{0}$, if $0\leq\mu\leq\mu_{0}$, no conflicts would occur with minimum aircraft separation requirements in $\bar{\phi}$ and if $\mu>\mu_{0}$, conflicts would occur with minimum aircraft separation requirements in $\bar{\phi}$. Then the constant $\mu_{0}$ is said to be the initial redundant time of the T-block subsequence $\phi$.
(2) Suppose that $Tcf_{i_{1}-1}\in\phi_{1}$, $Tcf_{i_{1}}\in\phi_{0}$, for the sequence $\phi_{1}$, there is a path from its last aircraft $Tcf_{i_{1}-1}$ to its first aircraft, and for the sequence $\phi_{0}$, there is a path from the second last aircraft to the first aircraft in $\phi_{0}$. If $Tcf_{i_{1}}$ is only relevant to $Tcf_{i_{1}-1}$, the sequence $\phi$ is a D-block subsequence of $\Phi$, it is said that $(|\phi_{0}|-1)/(|\phi_{1}|-1)$ and $S_{i_{1}}(\phi)-S_{i_{2}}(\phi)-Y_{i_{2}i_{1}}$ are the length and the landing time increment of D-block subsequence, where $Tcf_{i_{2}}$ is the second last landing aircraft in $\phi$. Let $\langle Tcf_{k_{0}},Tcf_{k_{1}},\cdots,Tcf_{k_{s}}\rangle$ denote the landing sequence in the D-block subsequence $\phi$. Consider a subsequence $\bar{\phi}$ such that the takeoff times of the takeoff aircraft in $\bar{\phi}$ are consistent with those in $\phi$ and $S_{k_{0}}(\bar{\phi})-S_{k_{0}}(\phi)=S_{k_{1}}(\bar{\phi})-S_{k_{1}}(\phi)=\cdots=S_{k_{s-1}}(\bar{\phi})-S_{k_{s-1}}(\phi)=\mu$ and $S_{k_{s}}(\bar{\phi})=S_{k_{s}}(\phi)$ for some constant $\mu\geq 0$. Suppose that for a given constant $\mu_{0}$, if $0\leq\mu\leq\mu_{0}$, no conflicts would occur with minimum aircraft separation requirements in $\bar{\phi}$ and if $\mu>\mu_{0}$, conflicts would occur with minimum aircraft separation requirements in $\bar{\phi}$. Then the constant $\mu_{0}$ is said to be the initial redundant time of the D-block subsequence $\phi$.
@@
In fact, under Assumption 17, the sequence $\Phi$ can be expressed as a group of block subsequences in the form of, e.g., $\langle TB_{1},TB_{2},DB_{1},TB_{3},DB_{2},DB_{3},TB_{4},\cdots\rangle$, where each $TB_{i}$ denotes a T-block subsequence and each $DB_{i}$ denotes a D-block subsequence. Based on the definitions of block subsequences, we can focus on the switching between T-block subsequences and D-block subsequences to study the time increments of the landing sequence and the takeoff sequence of the whole sequence $\Phi$.
@@
Moreover, it is assumed that $S_{1}(\phi_{0})=S_{1}(\phi_{1})$ in Tables I, III, V and VII and $S_{1}(\phi_{0})=S_{1}(\phi_{1})+60$ in Tables II, IV, VI and VIII.
@@
From Tables V-VIII, it can be seen that the range of the time increments of block subsequences is not large as well, in particular, when the takeoff time increments are considered. As a matter of fact, based on the calculations for more general scenarios including those listed in Tables I-VIII, it is found that the number of all possible block subsequences of different class combinations and the range of the takeoff/landing time increments of block subsequences are both not large according to the standards at the minimum separation time standards at Heathrow Airport and for the RECAT-EU system in the absence of resident-point aircraft. This property can be fully explored to find the optimal sequence when a landing sequence and a takeoff sequence are matched.

### v5

@@ VI Scheduling problem of landing and takeoff aircraft on dual runways
If one aircraft $Tcf_{i}$ lands and one aircraft $Tcf_{j}$ takes off at the same time, it is assumed by default that aircraft $Tcf_{i}$ is ahead of aircraft $Tcf_{j}$, i.e., $S_{i}\leq S_{j}$, the aircraft $Tcf_{i}$ is the leading aircraft and $Tcf_{j}$ is the trailing aircraft, and it is said that the aircraft $Tcf_{j}$ is relevant to the aircraft $Tcf_{i}$.
@@
Proof: For a landing aircraft $Tcf_{i}$ and a takeoff aircraft $Tcf_{k}$ in $\Phi$, if aircraft $Tcf_{k}$ is relevant to aircraft $Tcf_{i}$, then $S_{i}(\Phi)=S_{k}(\Phi)$. Note that $|S_{i}-S_{j}|\geq T_{0}$ for each $Tcf_{j}\in\Phi_{1}-\{Tcf_{i}\}$ and $|S_{k}-S_{h}|\geq T_{0}$ for each $Tcf_{h}\in\Phi_{0}-\{Tcf_{i}\}$, and $S_{h}\leq S_{j}$ for all $0<h<j$ from the sequence definition. Note that if one aircraft $Tcf_{a}$ lands and one aircraft $Tcf_{b}$ takes off at the same time, it is assumed by default that aircraft $Tcf_{a}$ is ahead of aircraft $Tcf_{b}$. Thus, $k=i+1$. For a takeoff aircraft $Tcf_{i}$ and a landing aircraft $Tcf_{k}$ in $\Phi$, if aircraft $Tcf_{k}$ is relevant to aircraft $Tcf_{i}$, $S_{k}(\Phi)-S_{i}(\Phi)=T_{0}$. Note that $|S_{i}-S_{j}|\geq T_{0}$ for each $Tcf_{j}\in\Phi_{1}-\{Tcf_{i}\}$ and $|S_{k}-S_{h}|\geq T_{0}$ for each $Tcf_{h}\in\Phi_{0}-\{Tcf_{i}\}$, and $S_{h}\leq S_{j}$ for all $0<h<j$ from the sequence definition. Also, note that if one aircraft $Tcf_{a}$ lands and one aircraft $Tcf_{b}$ takes off at the same time, it is assumed by default that aircraft $Tcf_{a}$ is ahead of aircraft $Tcf_{b}$. Since the takeoff aircraft $Tcf_{k}$ is relevant to the landing aircraft $Tcf_{i}$, then $k=i+1$.
@@
(2) Suppose that aircraft $Tcf_{i}$ and $Tcf_{i+1}$ are both landing (takeoff) aircraft. If aircraft $Tcf_{i+1}$ is relevant to some aircraft $Tcf_{k}$, then $k=i$.
@@
Proof: (1) From Lemma 16, this statement holds.
(2) From the sequence definition, $S_{h(i+1)}\geq S_{i(i+1)}\geq T_{0}$ for all $0<h<i$. If $k\neq i$, then $0<k<i$. Since $D_{P}=T_{0}$, if the landing aircraft $Tcf_{i+1}$ is relevant to a takeoff aircraft $Tcf_{k}$, then $Y_{k(i+1)}=T_{0}$. It is assumed that an aircraft lands and an aircraft takes off at the same time, it is assumed by default that the landing aircraft is ahead of the takeoff aircraft. Then the takeoff time of aircraft $Tcf_{k}$ is later than the landing time of aircraft $Tcf_{i}$, i.e., $k>i$. This yields a contradiction. Therefore, the landing aircraft $Tcf_{i+1}$ is relevant to a landing aircraft $Tcf_{k}$. From Theorem 1, $k=i$. If aircraft $Tcf_{i}$ and $Tcf_{i+1}$ are both takeoff aircraft, it can be proved that $k=i$.
(3) If one of the aircraft $Tcf_{i-1}$, $Tcf_{i+1}$ and $Tcf_{j-1}$, denoted by $Tcf_{h}$, is a landing (takeoff) aircraft, from Lemmas 1 and 7, $S_{ij}\geq Y_{ih}+Y_{hj}>Y_{ij}$, implying that aircraft $Tcf_{j}$ is not relevant to aircraft $Tcf_{i}$, which is a contradiction.
(4) From statements (1) and (3), this statement holds naturally.
(5) Since the aircraft $Tcf_{j}$ is relevant to aircraft $Tcf_{i}$, $Y_{ij}\leq 3T_{0}$. If aircraft $Tcf_{i}$ and $Tcf_{j}$ are both landing (takeoff) aircraft, $Tcf_{i+1},Tcf_{i+2},\cdots,Tcf_{j-1}$ are all takeoff (landing) aircraft. If $j-i>5$, it follows that $S_{ij}=S_{i(i+1)}+S_{(i+1)(j-1)}+S_{(j-1)j}>S_{(i+1)(j-1)}\geq 3T_{0}\geq Y_{ij}$, implying that aircraft $Tcf_{j}$ is not relevant to aircraft $Tcf_{i}$, which yields a contradiction. Suppose that one of aircraft $Tcf_{i}$ and $Tcf_{j}$ is a landing aircraft and one is a takeoff aircraft. If $j-i>5$, it is easy to see that $S_{ij}>T_{0}$, and aircraft $Tcf_{j}$ cannot be relevant to aircraft $Tcf_{i}$. Therefore, if the aircraft $Tcf_{j}$ is relevant to aircraft $Tcf_{i}$, $0<j-i\leq 5$.
@@
(2) If the optimization problem (6) can be solved, the aircraft $Tcf_{2}$ is relevant to the aircraft $Tcf_{1}$ and for each $i\in\{3,4,\cdots,n+m\}$, the aircraft $Tcf_{i}$ is relevant to an aircraft $Tcf_{k}$ for $k\in\{Tcf_{i-5},\cdots,Tcf_{i-2},Tcf_{i-1}\}$.
Proof: The proof is similar to that of Theorem 7.
@@
Proof: (1) In $\Phi$, $Tcf_{k+1}$ is located after $Tcf_{k}$. Thus, in $\Phi_{a}$, $Tcf_{k}$ is located after $Tcf_{k+1}$. From Remark 27 and Assumptions 16, 17, $S_{k}(\Phi_{a})=S_{k+1}(\Phi_{a})+T_{0}$. Since $Tcf_{k+1}$ is only relevant to $Tcf_{k}$ in $\Phi$, $S_{k+1}(\Phi)=S_{k}(\Phi)$. It follows that $S_{k+1}(\Phi)-S_{k+1}(\Phi_{a})+S_{k}(\Phi_{a})-S_{k}(\Phi)=T_{0}$.
(2) Similar to the proof of statement (1), this statement can be proved and the proof is omitted.
@@
In Theorem 13, the aircraft $Tcf_{k+1}$ is actually a semi-resident-point aircraft. From Theorem 13, the relative time increment is $T_{0}$ when the orders of aircraft $Tcf_{k}$ and $Tcf_{k+1}$ are exchanged, which is a significant property for the algorithm proposed later.
@@
Proof: From Theorem 13(1), $S_{i_{1}}(\Phi_{1})=S_{i_{0}}(\Phi_{0})$ and $S^{a}_{i_{0}}(\Phi_{0})-S^{a}_{i_{1}}(\Phi_{1})=T_{0}$, while from Theorem 13(2), $S_{i_{1}}(\Phi_{1})+T_{0}=S_{i_{0}}(\Phi_{0})$ and $S^{a}_{i_{0}}(\Phi_{0})=S^{a}_{i_{1}}(\Phi_{1})$. By simple calculations, this theorem can be proved.
@@
Proof: (1) To prove statement (1), we perform the analysis in two cases: (1.1) the aircraft $Tcf^{0}_{i+1}$ is not relevant to the aircraft $Tcf^{0}_{i}$ and (1.2) the aircraft $Tcf^{0}_{i+1}$ is relevant to the aircraft $Tcf^{0}_{i}$.
Case (1.1). Since the aircraft $Tcf^{0}_{i+1}$ is not relevant to the aircraft $Tcf^{0}_{i}$, $S_{j(j+3)}(\Phi_{1})\geq 3T_{0}$ and $D_{P}=T_{0}$, the aircraft $Tcf^{0}_{i+1}$ is relevant to the aircraft $Tcf^{1}_{k}$, $k\geq j+1$. If $k\geq j+3$, it can be checked that $S_{i(i+1)}(\Phi_{0})-T_{0}\geq Y_{i(i+1)}$, i.e., Assumption 17 does not hold. Thus, $k<j+3$. Note that $Y_{j(j+1)}(\Phi_{1})=Y_{(j+1)(j+2)}(\Phi_{1})=Y_{(j+2)(j+3)}(\Phi_{1})=T_{0}$. To ensure Assumption 17 to be valid, the aircraft should satisfy that $S_{(j+h)(j+h+1)}(\Phi_{1})=T_{0}$ for $h=0,1,\cdots,k$ and $S_{i+1}(\Phi_{0})-S_{k}(\Phi_{1})=T_{0}$. Clearly, $S_{i+1}(\Phi_{0})=S_{k+1}(\Phi_{1})$ where $k+1\in\{j+1,j+2,j+3\}$.
Case (1.2). Since $D_{P}=T_{0}$, for any integer $k\geq j$ if $S_{k}(\Phi_{1})\leq S_{i+1}(\Phi_{0})$, then $S_{i+1}(\Phi_{0})-S_{k}(\Phi_{1})\geq T_{0}$. Since the aircraft $Tcf^{0}_{i+1}$ is relevant to the aircraft $Tcf^{0}_{i}$, $T_{0}\leq S_{i(i+1)}(\Phi_{0})\leq 3T_{0}$. Thus, if $S_{k}(\Phi_{1})\leq S_{i+1}(\Phi_{0})$ for some integer $k\geq j$, under Assumption 17, it follows that $j\leq k\leq j+2$. Note that $S_{i+1}(\Phi_{0})-S_{j+k}(\Phi_{1})\geq T_{0}$ and $Y_{j(j+1)}(\Phi_{1})=Y_{(j+1)(j+2)}(\Phi_{1})=Y_{(j+2)(j+3)}(\Phi_{1})=T_{0}$. Thus, under Assumption 17, $S_{i+1}(\Phi_{0})=S_{k+1}(\Phi_{1})$ where $j\leq k\leq j+2$.
Summarizing the above analysis of both cases, there exists an integer $k_{0}\in\{j+1,j+2,j+3\}$ such that $S_{i+1}(\Phi_{0})=S_{k_{0}}(\Phi_{1})$.
(2)-(6) The proofs are similar to that for the statement (1) and hence omitted.
@@
From Theorem 16, it can be observed that the unequal separation times caused possibly by the existence of breakpoint aircraft and resident-point aircraft might result in mismatches between the landing sequence $\Phi_{0}$ and the takeoff sequence $\Phi_{1}$, but the effects of the unequal separation times would disappear after $1$ to $4$ aircraft in the absence of the influence of other breakpoint aircraft and resident-point aircraft. Moreover, it should be noted here that the forms of the minimum separation times with a sequence such as $Y_{j(j+1)}(\Phi_{1})$ and $Y_{(j+1)(j+2)}(\Phi_{1})$ are used instead of the forms $Y_{j(j+1)}$ and $Y_{(j+1)(j+2)}$ to distinguish different separation times between different sequences.
@@
(1) Suppose that $Tcf_{i_{1}-1}\in\phi_{0}$, $Tcf_{i_{1}}\in\phi_{1}$, for the sequence $\phi_{0}$, there is a path from its last aircraft $Tcf_{i_{1}-1}$ to its first aircraft, and for the sequence $\phi_{1}$, there is a path from its second last aircraft to its first aircraft. If $Tcf_{i_{1}}$ is relevant to only $Tcf_{i_{1}-1}$, it is said that the sequence $\phi$ is a T-block subsequence of $\Phi$, and $(|\phi_{1}|-1)/(|\phi_{0}|-1)$ and $S_{i_{1}}(\phi)-S_{i_{2}}(\phi)-Y_{i_{2}i_{1}}$ are the length and the takeoff time increment of T-block subsequence, where $Tcf_{i_{2}}$ is the second last takeoff aircraft in $\phi$. Let $\langle Tcf_{k_{0}},Tcf_{k_{1}},\cdots,Tcf_{k_{s}}\rangle$ denote the takeoff sequence in the T-block subsequence $\phi$. Consider a subsequence $\bar{\phi}$ such that the landing times of the landing aircraft in $\bar{\phi}$ are consistent with those in $\phi$ and the takeoff times of the takeoff aircraft satisfy that $S_{k_{0}}(\bar{\phi})-S_{k_{0}}(\phi)=S_{k_{1}}(\bar{\phi})-S_{k_{1}}(\phi)=\cdots=S_{k_{s-1}}(\bar{\phi})-S_{k_{s-1}}(\phi)=\mu$ and $S_{k_{s}}(\bar{\phi})=S_{k_{s}}(\phi)$ for some constant $\mu\geq 0$. Suppose that for a given constant $\mu_{0}$, if $0\leq\mu\leq\mu_{0}$, no conflicts of separation times would occur in $\bar{\phi}$ and if $\mu>\mu_{0}$, conflicts of separation times would occur in $\bar{\phi}$. Then the constant $\mu_{0}$ is said to be the initial redundant time of the T-block subsequence $\phi$.
(2) Suppose that $Tcf_{i_{1}-1}\in\phi_{1}$, $Tcf_{i_{1}}\in\phi_{0}$, for the sequence $\phi_{1}$, there is a path from its last aircraft $Tcf_{i_{1}-1}$ to its first aircraft, and for the sequence $\phi_{0}$, there is a path from the second last aircraft to the first aircraft in $\phi_{0}$. If $Tcf_{i_{1}}$ is relevant to only $Tcf_{i_{1}-1}$, it is said that the sequence $\phi$ is a D-block subsequence of $\Phi$, and $(|\phi_{0}|-1)/(|\phi_{1}|-1)$ and $S_{i_{1}}(\phi)-S_{i_{2}}(\phi)-Y_{i_{2}i_{1}}$ are the length and the landing time increment of D-block subsequence, where $Tcf_{i_{2}}$ is the second last landing aircraft in $\phi$. Let $\langle Tcf_{k_{0}},Tcf_{k_{1}},\cdots,Tcf_{k_{s}}\rangle$ denote the landing sequence in the D-block subsequence $\phi$. Consider a subsequence $\bar{\phi}$ such that the takeoff times of the takeoff aircraft in $\bar{\phi}$ are consistent with those in $\phi$ and $S_{k_{0}}(\bar{\phi})-S_{k_{0}}(\phi)=S_{k_{1}}(\bar{\phi})-S_{k_{1}}(\phi)=\cdots=S_{k_{s-1}}(\bar{\phi})-S_{k_{s-1}}(\phi)=\mu$ and $S_{k_{s}}(\bar{\phi})=S_{k_{s}}(\phi)$ for some constant $\mu\geq 0$. Suppose that for a given constant $\mu_{0}$, if $0\leq\mu\leq\mu_{0}$, no conflicts of separation times would occur in $\bar{\phi}$ and if $\mu>\mu_{0}$, conflicts of separation times would occur in $\bar{\phi}$. Then the constant $\mu_{0}$ is said to be the initial redundant time of the D-block subsequence $\phi$.
@@
In fact, under Assumption 17, the sequence $\Phi$ can be expressed as a sequence of block subsequences in the form of, e.g., $\langle TB_{1},TB_{2},DB_{1},TB_{3},DB_{2},DB_{3},TB_{4},\cdots\rangle$, where each $TB_{i}$ denotes a T-block subsequence and each $DB_{i}$ denotes a D-block subsequence. It should be noted that for any two consecutive block subsequences, the last two aircraft of the first block subsequence are the first two aircraft of the second block subsequences. Based on the definitions of block subsequences, we can focus on the switching between T-block subsequences and D-block subsequences to study the time increments of the landing sequence and the takeoff sequence of the whole sequence $\Phi$.
According to the combination of the first two aircraft, each T-block/D-block subsequence has two different forms. One form is that the first aircraft is a landing aircraft, the second aircraft is a takeoff aircraft and the separation time between these two aircraft is 00. The other form is that the first aircraft is a takeoff aircraft, the second aircraft is a landing aircraft and the separation time between these two aircraft is $T_{0}$.
@@
Moreover, it is assumed that $S_{1}(\phi_{0})=S_{1}(\phi_{1})$ in Tables I, III, V and VII, and $S_{1}(\phi_{0})=S_{1}(\phi_{1})+60$ in Tables II, IV, VI and VIII.
@@
From Tables V-VIII, it can be seen that the range of the time increments of block subsequences is not large as well, in particular, when the takeoff time increments are considered. As a matter of fact, based on the calculations for more general scenarios including those listed in Tables I-VIII, it is found that if the number of breakpoint aircraft in a block subsequence is fixed, the number of all possible block subsequences of different class combinations and the range of the takeoff/landing time increments of block subsequences are both not large according to the minimum separation time standards at Heathrow Airport and for the RECAT-EU system in the absence of resident-point aircraft. This property can be fully explored to find the optimal sequence when a landing sequence and a takeoff sequence are matched.